OpenStudy (anonymous):

let A be 3x3 matrix with (2,0,-1) as first row (4,0,-2) as second row and (0,0,0) as third row .let T be multiplication by A find basis for range of T

OpenStudy (anonymous):

2 0 -1 4 0 -2 0 0 0 ?

OpenStudy (anonymous):

yes this is my 3x 3 matrix

OpenStudy (anonymous):

o -2 4 -2 4 0 (2) 0 0 (-0) 0 0 +(-1) 0 0

OpenStudy (anonymous):

sir, please explain i am not getting your vpoint

OpenStudy (anonymous):

I,m not done yet

OpenStudy (anonymous):

read this is hard for explain

OpenStudy (anonymous):

[(2)(0*0)-(0*-2) (-0)(0*4)-(0*-2) (-1)(0*4)-(0*0)] [(2*2) ( ............... time 0 all cancel..........] =4

OpenStudy (anonymous):

first use sign + - +

OpenStudy (anonymous):

the first row 2 0 -1 , (2 is positive sign) ,( 0 negative sign) , ( -1 positive sign)

OpenStudy (anonymous):

if you are trying to find a basis for the range of the transformation, you should row reduce the matrix, and find the pivots. those pivots indicate which columns are linearly independent and form a basis for the range of T. I dont know of anyway to use the determinant of a matrix to find a basis for the range.....although i wont claim there isnt a way. i just dont know of it if there is.

OpenStudy (anonymous):

(2) minor is : 0 -2 0 0 you cross (0*0)-(0*-2)

OpenStudy (anonymous):

0-(-2)=2 2*2=4

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

(-0) minor is : 4 -2 0 0 you cross (4*0)-(0*-2) (-0)* 0 -(-2) = -0*2 = 0

OpenStudy (anonymous):

4 0 +(-1) 0 0 you cross and time= (-1)* (4*0)-(0*0) ( -1*) 0 -0 = -1*0 = 0

OpenStudy (anonymous):

the answer is 4

OpenStudy (anonymous):

or you can solve by use Diagonals,

OpenStudy (anonymous):

not hard to solve, but I don't know how to explain on the chat, my son love to solve Matrices

OpenStudy (anonymous):

i dont mean to sound rude, but I dont see how this solves his problem at all. His problem was asking for a basis for the range of T, not the determinant of A.

OpenStudy (anonymous):

than I don't know

OpenStudy (anonymous):

i posted this further up, here is the solution:

OpenStudy (anonymous):

@sir joemath the answer in my book is (1,2,0) i got your point very well but the answer in book is different from yours

OpenStudy (anonymous):

ah, we have the same answer, their answer is mine divided by 2. the concept is like reducing fractions: \[\frac{2}{4} = \frac{1}{2}\] i guess they want to pull out any common factors between the numbers in their basis? i dont see why you would want to do that, but hey, when in Rome.

OpenStudy (anonymous):

@ sir joemaththanks thanks thanks sir for answer nancy lam was telling me about determinent but i was asking about range which you told me

OpenStudy (anonymous):

[(2)(0*0)-(0*-2) (-0)(0*4)-(0*-2) (-1)(0*4)-(0*0)] [(2*0) ( ............... time 0 all cancel..........] =0 I calculator wrong is o

OpenStudy (anonymous):

it was just a misunderstanding, it happen to the best of us, ive done it a bajillion times lol

OpenStudy (anonymous):

The basis for the range is the same as the basis for the column space if I'm not mistaken. . .

OpenStudy (anonymous):

thats correct :)